OpenStudy (anonymous):

Write an equation and determine the 8th term: 7, 7.5, 7.75, 7.875

OpenStudy (anonymous):

All I really need to know is the common difference :s

OpenStudy (isaiah.feynman):

The difference increases by a factor of 0.5.

OpenStudy (jdoe0001):

hmmm.... 7.5 to 7.75 is just 0.25

OpenStudy (isaiah.feynman):

But 7 to 7.5 is 0.5. :P

OpenStudy (jdoe0001):

heheh =)

OpenStudy (jdoe0001):

actually I noticed the ratio is an exponential anyhow \(\begin{array}{lllll} &7+0.5^{1-1}&7+0.5^{2-1}&7+0.5^{3-1}&7+0.5^{4-1}\\ \hline\\ \implies &7& 7.5& 7.75& 7.875\\ \hline\\ n^{th}&1&2&3&4 \end{array}\)

OpenStudy (agent0smith):

There's no common difference. But that pizza looks delicious. @jdoe0001 it looks more like a series sum. 7+0.5+0.5^2+0.5^3...

OpenStudy (jdoe0001):

well... hmm... well... sorta, yes

OpenStudy (agent0smith):

The next term is found by adding 0.5^(n-1)... Oh but it starts with 6, \[\Large 6+0.5^0+0.5^1+0.5^2+0.5^3...\]

OpenStudy (jdoe0001):

hmmm... ohh right... the 1st item.... will be 8 otherwise.

OpenStudy (jdoe0001):

or one can say it starts with 7, and sequences out to the 2nd and forth

OpenStudy (jdoe0001):

hmmm wait... nope you're right.. it starts with 6 either way

OpenStudy (jdoe0001):

hmmm ahemm....

OpenStudy (jdoe0001):

\(\bf \begin{array}{ccccccc} &&\color{red}{(n-1)+0.5^{1-1}}&7+0.5^{2-1}&7.5+0.5^{3-1}&7.75+0.5^{4-1}\\ \hline\\ \implies &6&7& 7.5& 7.75& 7.875\\ \hline\\ n^{th}&1&2&3&4&5 \end{array}\)

OpenStudy (jdoe0001):

woops.. forgot something

OpenStudy (agent0smith):

You can use a geometric sequence sum to find it, just add the 6 at the end... each term is just 6 plus the geometric sum to that point ignoring the 6, the geometric sequence part is \[\Large a_n = a_1 r^{n-1}\]and the sum for a geometric is \[\Large a_1 \left( \frac{ 1-r^n }{ 1-r } \right)\] your first term is 1, common ratio is 0.5. \[\Large a_n = 0.5^{n-1}\]

OpenStudy (jdoe0001):

he's offline right now, so no pizza for you yet =)

OpenStudy (agent0smith):

So the 8th term \[\Large 6+ \Large \left( \frac{ 1-0.5^8 }{ 1-8 } \right)\]

OpenStudy (agent0smith):

I guess a general formula would be \[\Large a_n = 6 + \Large \left( \frac{ 1-0.5^n }{ 1-0.5 } \right)\]

OpenStudy (agent0smith):

I think that works for all the sequence... never really needed to work out a sequence of this type tho, so the explanation isn't very clear as i was figuring it out as i went along.

OpenStudy (jdoe0001):

hmm if I take the 7 as the 1st term.... \(\bf (n-1)+0.5^{n-1}\) seem to work out as the common ratio

OpenStudy (agent0smith):

Yeah i guess that would would too. But the above formula works and can be simplified to \[\Large a_n = 8-2*0.5^n \]

OpenStudy (agent0smith):

Looks better this way \[\Large a_n = 8-2(0.5^n)\]

OpenStudy (anonymous):

@agent0smith @jdoe0001 oh my, seems complicated, but I think i got it down! thanks for the help :)